JoelTheBoss (joel_the_boss):

Let f(x) = 3x2 + x - 3 and g(x) = x2 - 5x + 1. Find f(x) - g(x). 2x2 - 4x - 2 2x2 - 4x - 4 2x2 + 6x - 2 2x2 + 6x - 4

JoelTheBoss (joel_the_boss):

@iGreen @ganeshie8

JoelTheBoss (joel_the_boss):

@undeadknight26

OpenStudy (anonymous):

Let f(x) = 3x2 + x - 3 and g(x) = x2 - 5x + 1. Find f(x) - g(x). okay so what you do is this. you have the equations, they want you to subtract them. But I'm guessing you mean 3x^2 right? and x^2

JoelTheBoss (joel_the_boss):

yes I do

OpenStudy (anonymous):

So if that is so: 3x^2 + x-3 - x^2-5x+1 combine similar terms you get 2x^2 -4x - 2

OpenStudy (yanasidlinskiy):

f(x) – g(x) f(x) = 3x^2 + x – 3 g(x) = x^2 – 5x + 1 so 3x^2 + x – 3 - (x^2 – 5x + 1) 3x^2 + x - 3 - x^2 +5x -1 Combine common factors 3x^2 - x^2 x + 5x - 3 - 1 And you get: 2x^2 + 6x - 4

OpenStudy (anonymous):

lol yana how did you get that?

OpenStudy (yanasidlinskiy):

Lol, sorry!! I hate those things. I don't know why it turned out to be that way!! Sorry again!!

OpenStudy (anonymous):

ahh you multiplied by -1

OpenStudy (yanasidlinskiy):

Lol, had to refresh!!xD